For the system of linear equation
-Determine how many solutions exist
-Use either elimination or substitution to find the solutions (if any)
-Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection
x + y = 3, y = x + 3
Here is what I have done so far using substitution:
x + y = 3, y = x + 3
x + y = 3
x + y – y = 3 – y
x = 3 – y
3 – y + y = 3 – y
3 = 3 – y
3 – 3 = 3 – 3 – y
y = 0
x + 0 = 3
x = 3
check
x + y = 3
3 + 0 = 3
3 = 3 (It checks here but not in the second equation)
It seems I am getting the reverse of what the answer should be. I have done it many times and I get the same answer. Please help!

Answer by Alan3354(69443)

x + y = 3, y = x + 3
---------
x + y = 3
x - y = -3
------------- Add
2x = 0
x = 0
----
y = 3